# Which Reasoning Trajectories Teach Students to Reason Better? A Simple Metric of Informative Alignment

### v3

@@ 1 Introduction
To address this challenge, we propose a simple yet effective metric, Rank-Surprisal Ratio (RSR), which quantifies the suitability of a reasoning trajectory for a given student by jointly capturing alignment and informativeness. Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting both absolute unfamiliarity and relative familiarity. Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while still containing behavioral patterns within the student’s prior experience, such that their tokens rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
Figure 1: Illustration of the intuition behind RSR. Suitable reasoning trajectories should balance informativeness and alignment, having low absolute probability while their tokens remain relatively high-ranked under the student model.
@@
  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data-student suitability.
  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory for a given student model by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
  • We demonstrate the practical utility of Rank-Surprisal Ratio in two data engineering scenarios, trajectory selection and teacher selection, where it serves as an effective criterion and outperforms existing methods.
@@ 2 The Need for Student-Specific Data
To understand which types of reasoning trajectories most effectively improve student models after SFT, we conduct a comprehensive large-scale study involving five widely adopted student models and eleven diverse reasoning-oriented teacher models, yielding 55 teacher–student pairings. We perform SFT experiments for each of these pairs.
@@
Table 1: Distillation results showing post-training reasoning performance of student models trained on trajectories from different teacher models, evaluated by average Acc@4 on AIME’25, AIME’24, AMC’23, and MATH500. Darker and lighter shading indicate the best and second-best results, respectively. Student performance varies significantly across teacher–student pairs, highlighting the importance of data–student suitability.
@@ 2.1 Experimental Settings
Our teacher-student pairing study involves two major steps: (1) For each teacher model, we prompt it to generate a long CoT response for each math problem in our 5000-problem set (see § A.1), forming a trajectory dataset specific to that teacher. (2) For each teacher–student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance. All students are pre-trained base models. To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher–student pair. Reported results are averaged over these three runs. More implementation details are provided in Appendix A.
@@ 2.2 Results
Table 1 presents the results of our teacher–student pairing distillation study, revealing that:
@@
Data–student suitability is critical for eliciting reasoning improvements.
The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model. Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) can likewise be ineffective when training stronger students (Qwen-3-14B). Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb. Overall, we find no simple teacher–student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
@@ 3 Measuring Data-Student Suitability
In this section, we explore metrics for measuring data–student suitability, with the goal of jointly capturing informativeness and alignment. We begin by introducing two fundamental token-level measures: surprisal and rank (§ 3.1), and analyzing the limitations of existing probability-based metrics (§ 3.2). We then abstract our insights on suitable reasoning data and conduct simulation studies to identify quantitative patterns (§ 3.3). Finally, we propose our trajectory-level metric (§ 3.4).
@@
Table 2: Comparison of the student’s post-training performance and data–student suitability metrics across trajectories from different teacher models, evaluated on Qwen-2.5-7B. Darker shading indicates higher performance or better suitability. Metrics whose trends align with performance (e.g., RSR) provide more reliable suitability estimates. Complete metric scores are provided in § C.2.
@@ 3.3 Insight and Simulation
Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while still exhibiting behavioral patterns within the student’s prior experience. Therefore, tokens in such trajectories are assigned low absolute probability (i.e., high surprisal) by the student model while still ranking relatively high (i.e., having low rank values) in its prediction distribution.
@@
We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we deem effective. Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$. The first mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant patterns, which arise from student’s massive general training data. The second mode $Z_{B}$, represents tokens whose patterns resemble strong reasoning trajectories, reflecting the student’s prior exposure to relevant data. We instantiate both $Z_{A}$ and $Z_{B}$ as Zipf distributions [54, 27] over $\mathcal{V}$. Then the student’s overall token-level prediction distribution $Z$ is constructed as a mixture of $Z_{A}$ and $Z_{B}$:
@@
Based on $Z$, we simulate four types of reasoning trajectories and examine their average token surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior. More simulation details are provided in § A.6.
Figure 2: Simulation of the student model’s token-level prediction distribution $Z$. Tokens with low probability but relatively high rank—characteristic of the $Z_{B}$ mode—yield smaller token-level rank–surprisal ratios.
@@
$X_{A}$ (from $Z_{A}$) 0.41 1.38 2.49 1.69
$X_{B}$ (from $Z_{B}$) 0.10 2.73 4.31 1.30
$X_{C}$ (from $Z_{C}$) 0.08 4.73 11.57 2.23
$X_{D}$ (from $Z$) 0.35 1.67 2.93 1.62
@@ 3.4 Proposed Metric: Rank-Surprisal Ratio
A natural solution is to exclude tokens with very low surprisal when computing the average. Let $\mathcal{T}_{H}(\mathbf{x})$ denote the set of response tokens whose surprisal lies in the top $H\%$ within a trajectory $\mathbf{x}$. We define a filtered average as
@@
Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2). This suggests that tokens with higher surprisal have a greater impact on student learning and should therefore be emphasized when computing the average.
@@
The resulting metric yields a concise form that can be interpreted as the ratio of the average token rank to the average token surprisal over a trajectory.
In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability. Since tokens with excessively large ranks are effectively indistinguishable from the student’s perspective, we clip rank values at a threshold $r_{max}$. Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR [^2] By default, RSR refers to the trajectory-level RSR in this paper, unless otherwise specified (e.g., in correlation analysis).), as
@@
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates greater deviation from dominant patterns and provides more informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision. § C.5 presents illustrative examples of RSR measurements.
@@ 4.1 Main Analysis
Table 4: Spearman correlation between different metrics and post-training reasoning performance across student models, reporting absolute values. "Student-Agnostic" metrics are computed independently of the specific student model. Our metric, Rank-Surprisal Ratio, achieves the highest correlation across all students.
For each of the five student models and each metric, we aggregate trajectory-level suitability (or quality) scores to obtain dataset-level scores for reasoning datasets generated by eleven teacher models (§ 2.1). We then measure the correlation between these dataset-level scores and the student’s reasoning performance after training on the corresponding teacher-generated datasets. We primarily report Spearman’s correlation coefficient, while Pearson correlation exhibits similar trends (§ C.2). For dataset-level RSR, we adopt a weighted averaging scheme (§ A.8) similar to Eq. 6, which yields slightly higher correlation than a simple average of trajectory-level RSR. We use a clipping threshold of $r_{max}=100$ for RSR in all subsequent experiments. Additional analysis details are provided in § A.7, and complete metric scores are reported in § C.2.
@@
We compare RSR against a diverse set of metrics for evaluating reasoning trajectories. These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal [19]) and rank-based metrics. We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [45], LLM-judged quality scores, and answer accuracy on verifiable questions. In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE [30]) and influence scores [17].
@@ Results
Table 4 shows that Rank-Surprisal Ratio consistently exhibits strong correlation with post-training reasoning performance across all student models, achieving an average Spearman correlation of 0.86 and outperforming all alternative metrics. These results indicate the effectiveness and practical value of RSR. In contrast, surprisal-based and rank-based metrics alone yield substantially weaker correlations (at most 0.59), highlighting the importance of capturing both informativeness and alignment through the rank-surprisal ratio.
@@ 4.2 Ablation Study
As shown in Table 5, removing either rank clipping or the surprisal-weighted averaging substantially degrades the correlation, validating the necessity of both components in our metric. In addition, the “Reduced sample size” setting estimates the dataset-level RSR using only 200 trajectories per teacher instead of the full 5,000. The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$. Additional ablation results are in § C.3.
@@ 5.1 Trajectory Selection
Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods (5k samples each). maxmax and minmin indicate maximizing or minimizing the corresponding metric. Model names are abbreviated as Q for Qwen and L for LLaMA. Additional results, including the GPQA evaluation and the No-Selection baseline, are provided in § C.4.
@@ Results
Additional trajectory selection results, including complete scores, further analysis, the No-Selection baseline, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
@@ 5.2 Teacher Selection
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation, reflecting a recurring challenge in real-world applications. To better capture practical constraints, we consider a low-resource setting in which generating full training data for every candidate teacher model is either costly or infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, evaluate them using different metrics, and select the teacher model based on the resulting dataset-level average score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@ 6 Related Work
Knowledge distillation is a powerful approach for transferring knowledge from large models to smaller ones [16], and has been widely used in training LLMs [31]. Prior work shows that stronger teachers do not necessarily yield better students, often due to capability gaps [48, 25] or off-policy data [5]. Recent studies address this by better aligning teacher supervision with student behavior, achieving an implicit balance through approaches such as on-policy distillation [2], integration with reinforcement learning [26, 51], SFT optimization [44, 39], teaching assistants [28, 9], and interleaved sampling [40, 32]. By contrast, our work explicitly quantifies the trade-off between informativeness and alignment in distillation, enabling a principled identification of effective teacher data for a given student.
@@
Long CoT trajectories provide strong supervision for improving student models’ reasoning performance via SFT [37, 35]. In line with findings in the general domain that high-quality data improves SFT [53, 43], many studies focus on constructing high-quality CoT data, either by selecting better prompts [46, 42] or by filtering reasoning trajectories [45, 18, 38, 55, 6, 24]. Recognizing that optimal reasoning data may vary across students, recent work explores student-specific trajectory selection strategies [30, 19, 17]. Our work also studies student-specific data selection, but differs from prior work by measuring data–student suitability from the perspective of informative alignment and by conducting more comprehensive evaluations across a wider range of teacher models.
7 Conclusion
In this paper, we study data–student suitability in reasoning distillation and propose Rank-Surprisal Ratio (RSR), a simple metric for identifying suitable reasoning trajectories for a given student. Motivated by our analysis, RSR jointly captures a trajectory’s informativeness and alignment with the student’s behavior, favoring trajectories with low absolute probability but relatively high-ranked tokens. Experiments across diverse teacher–student pairs show that RSR strongly correlates with post-training performance and consistently outperforms existing metrics. We further demonstrate its effectiveness in both trajectory and teacher selection. Overall, our results highlight informative alignment as a promising direction for reasoning distillation.

### v4

@@ 1 Introduction
To address this challenge, we propose a simple yet effective metric, Rank-Surprisal Ratio (RSR), which quantifies the suitability of a reasoning trajectory for a given student by jointly capturing alignment and informativeness. Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting both absolute unfamiliarity and relative familiarity. Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while remaining compatible with its overall generation patterns, such that their tokens still rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
Figure 1: Illustration of the intuition behind Rank-Surprisal Ratio. Suitable reasoning trajectories should balance informativeness and alignment, having low absolute probability while their tokens remain relatively high-ranked under the student model.
@@
  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data-student suitability (§ 2).
  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory for a given student model by jointly capturing alignment and informativeness (§ 3), achieving a strong correlation with post-training performance (§ 4).
  • We demonstrate the practical utility of RSR in two data engineering scenarios, trajectory selection and teacher selection, where it serves as an effective criterion and outperforms existing methods (§ 5). We further discuss the broader applicability of RSR to non-CoT data and subset selection (§ 7).
@@ 2 The Need for Student-Specific Data
To understand which types of reasoning trajectories most effectively improve student models after SFT, we conduct a comprehensive large-scale study involving five widely adopted student models and eleven diverse reasoning-oriented teacher models, yielding 55 teacher-student pairings. We perform SFT experiments for each of these pairs.
@@
Table 1: Distillation results showing post-training reasoning performance of student models trained on trajectories from different teacher models, evaluated by average Acc@4 on AIME’25, AIME’24, AMC’23, and MATH500. Darker and lighter shading indicate the best and second-best results, respectively. Student performance varies significantly across teacher-student pairs, highlighting the importance of data-student suitability.
@@ 2.1 Experimental Settings
Our teacher-student pairing study involves two major steps: (1) For each teacher model, we prompt it to generate a long CoT response for each math problem in our 5000-problem set (see § A.1), forming a trajectory dataset specific to that teacher. (2) For each teacher-student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance. All students are pre-trained base models. To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher-student pair. Reported results are averaged over these three runs. More implementation details are provided in Appendix A.
@@ 2.2 Results
Table 1 presents the results of our teacher-student pairing distillation study, revealing that:
@@
Data-student suitability is critical for eliciting reasoning improvements.
The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model. Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) can likewise be ineffective when training stronger students (Qwen-3-14B). Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb. Overall, we find no simple teacher-student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
@@ 3 Measuring Data-Student Suitability
In this section, we explore metrics for measuring data-student suitability, with the goal of jointly capturing informativeness and alignment. We begin by introducing two fundamental token-level measures: surprisal and rank (§ 3.1), and analyzing the limitations of existing probability-based metrics (§ 3.2). We then abstract our insights on suitable reasoning data and conduct simulation studies to identify quantitative patterns (§ 3.3). Finally, we propose our trajectory-level metric (§ 3.4).
@@
Table 2: Comparison of the student’s post-training performance and data-student suitability metrics across trajectories from different teacher models, evaluated on Qwen-2.5-7B. Darker shading indicates higher performance or better suitability. Metrics whose trends align with performance (e.g., RSR) provide more reliable suitability estimates. Complete metric scores for all teacher and student models are provided in § F.
@@ 3.3 Insight and Simulation
Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while remaining compatible with the student’s overall generation patterns learned from prior experience. Therefore, tokens in such trajectories are assigned low absolute probability (i.e., high surprisal) by the student model while still ranking relatively high (i.e., having low rank values) in its prediction distribution.
@@
We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we deem effective. Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$. The first (major) mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant generation patterns, which arise from its general training data. The second (minor) mode, $Z_{B}$, occupies a small fraction of the probability mass. It represents tokens from specific generation patterns that deviate from the major mode yet remain familiar to the student model. We instantiate both $Z_{A}$ and $Z_{B}$ as Zipf distributions [56, 27] over $\mathcal{V}$. Then the student’s overall token-level prediction distribution $Z$ is constructed as a mixture of $Z_{A}$ and $Z_{B}$:
@@
Based on $Z$, we simulate four types of reasoning trajectories and examine their average token surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant generation patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior. More simulation details are provided in § A.6.
Figure 2: Simulation of the student model’s token-level bimodal prediction distribution $Z$. Tokens with low probability yet ranked near the top (typical of the minor mode $Z_{B}$) tend to yield smaller, and thus preferred, rank-surprisal ratios.
@@
$X_{A}$ (from $Z_{A}$, major) 0.41 1.38 2.49 1.69
$X_{B}$ (from $Z_{B}$, minor) 0.10 2.73 4.31 1.30
$X_{C}$ (from $Z_{C}$, misaligned) 0.08 4.73 11.57 2.23
$X_{D}$ (from $Z$, mixed) 0.35 1.67 2.93 1.62
@@ 3.4 Proposed Metric: Rank-Surprisal Ratio
A natural solution is to exclude tokens with very low surprisal when computing the average. Let $\mathcal{T}_{H}(\mathbf{x})$ denote the set of tokens whose surprisal lies in the top $H\%$ within a trajectory $\mathbf{x}$. We define a filtered average as
@@
Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2). This suggests that response tokens with higher surprisal have a greater impact on student learning and should therefore be emphasized when computing the average.
@@
The resulting metric yields a concise form, interpretable as the ratio of the average token rank to the average token surprisal over a trajectory’s response tokens.
In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability. Since such tokens are effectively indistinguishable to the student, we clip rank values at a threshold $r_{max}$ to improve stability without sacrificing meaningful information. Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR) [^2] By default, RSR refers to the trajectory-level RSR in this paper, unless otherwise specified (e.g., in correlation analysis)., as
@@
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates greater deviation from dominant patterns and provides more informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision. § D.2 presents illustrative examples of RSR measurements.
@@ 4.1 Main Analysis
Table 4: Spearman correlation between each data-student suitability metric and post-training reasoning performance (average accuracy on reasoning benchmarks), reported in absolute values for different student models. “Student-agnostic” metrics are computed independently of the specific student model.
For each of the five student models and each metric, we aggregate trajectory-level suitability (or quality) scores to obtain dataset-level scores for reasoning datasets generated by eleven teacher models (§ 2.1). We then measure the correlation between these dataset-level scores and the student’s reasoning performance after training on the corresponding teacher-generated datasets. We primarily report Spearman’s correlation coefficient, while Pearson correlation exhibits similar trends (§ E.1). For dataset-level RSR, we adopt a weighted averaging scheme (§ A.8) similar to Eq. 6, which yields slightly higher correlation than a simple average of trajectory-level RSR. We use a clipping threshold of $r_{max}=100$ for RSR in all subsequent experiments. Additional analysis details are provided in § A.7, and complete metric scores are reported in § F.
@@
We compare RSR against a diverse set of metrics for evaluating reasoning trajectories. These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal [19]) and rank-based metrics. We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [46], LLM-judged quality scores, and answer accuracy on verifiable questions. In addition, we include other recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE [30]) and influence scores [17].
@@ Results
Table 4 shows that RSR consistently exhibits strong correlation with post-training reasoning performance across all student models, achieving an average Spearman correlation of 0.86 and outperforming all alternative metrics. These results indicate the effectiveness and practical value of RSR. In contrast, surprisal-based and rank-based metrics alone yield substantially weaker correlations (at most 0.59). Analysis (e.g., § 3.2) suggests that they tend to emphasize high-likelihood trajectories while insufficiently capturing informativeness. This highlights the importance of jointly modeling informativeness and alignment via the rank-surprisal ratio.
@@ 4.2 Ablation Study
As shown in Table 5, removing either rank clipping or the surprisal-weighted averaging substantially degrades the correlation, validating the necessity of both components in our metric. In addition, the “Reduced sample size” setting estimates the dataset-level RSR using only 200 trajectories per teacher instead of the full 5,000. The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$. Additional ablation results are in § E.2.
@@ 5.1 Trajectory Selection
Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods (5k samples each). maxmax and minmin indicate maximizing or minimizing the corresponding metric. Model names are abbreviated as Q for Qwen and L for LLaMA. Further results, including complete scores and additional GPQA evaluation, are provided in § F and § E.3.
@@ Results
Complete scores are reported in § F. Additional trajectory selection results, including GPQA evaluation, No-Selection baseline, further analyses, and reduced-teacher experiments, are provided in § E.3.
@@ 5.2 Teacher Selection
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation, reflecting a recurring challenge in real-world applications. To better capture practical constraints, we consider a low-resource setting in which generating full training data for every candidate teacher model is either costly or infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, evaluate them using data-student suitability metrics, and select the teacher model based on the resulting dataset-level average score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@ 6 Related Work
Knowledge distillation is a powerful approach for transferring knowledge from large models to smaller ones [16], and has been widely used in training LLMs [31]. Prior work shows that stronger teachers do not necessarily yield better students, often due to capability gaps [50, 25] or off-policy data [5]. Recent studies address this by better aligning teacher supervision with student behavior, achieving an implicit balance through approaches such as on-policy distillation [2], self-distillation [49, 35], integration with reinforcement learning [26, 53], SFT optimization [45, 40], teaching assistants [28, 9], uncertainty-based filtering [51], and interleaved sampling [41, 32]. By contrast, our work explicitly quantifies the trade-off between informativeness and alignment in distillation, enabling a principled identification of effective teacher data for a given student.
@@
Long CoT trajectories provide strong supervision for improving student models’ reasoning performance via SFT [38, 36]. In line with findings in the general domain that high-quality data improves SFT [55, 44], many studies focus on constructing high-quality CoT data, either by selecting better prompts [47, 43] or by filtering reasoning trajectories [46, 18, 39, 57, 6, 24]. Recognizing that optimal reasoning data may vary across students, recent work explores student-specific trajectory selection strategies [30, 19, 17]. Our work also studies student-specific data selection, but differs from prior work by measuring data-student suitability from the perspective of informative alignment and by conducting more comprehensive evaluations across a wider range of teacher models.